# Gradient Flow Finite Element Discretisations with Energy-Based $hp$–Adaptivity for the Gross-Pitaevskii Equation with Angular Momentum Rotation

### v1

@@ 3.1. $hp$-Finite element discretisation
For the purpose of a computational realisation of the discrete GFI (38), the Galerkin spaces $\mathbb{X}_{N}$ will be constructed in terms of an $hp$–finite element approach. We focus on a sequence of conforming and shape-regular partitions $\{\mathcal{T}_{N}\}_{N\in\mathbb{N}}$ of the domain $\Omega$ into simplicial elements $\mathcal{T}_{N}=\{\kappa\}_{\kappa\in\mathcal{T}_{N}}$ (i.e., triangles for $d=2$ and tetrahedra for $d=3$), although more general elements could also be considered. For $N\geq 1$ and an element $\kappa\in\mathcal{T}_{N}$, we denote by $h_{\kappa}:=\mbox{diam}(\kappa)$ the diameter of $\kappa$. Furthermore, to each element $\kappa\in\mathcal{T}_{N}$, we associate a polynomial degree $p_{\kappa}$, $p_{\kappa}\geq 1$, and collect these quantities in the polynomial degree vector $\bm{\mathsf{p}}_{N}=[p_{\kappa}:\kappa\in\mathcal{T}_{N}]$, $N\geq 1$. With this notation, for any subset $\varpi\subset\mathcal{T}_{N}$, we introduce the following $hp$–finite element space:
@@
$\int_{\Omega}\left(\frac{1}{2}\nabla u_{N}\cdot\nabla v+V(\bm{\mathsf{x}})u_{N}v+\beta|u_{N}|^{2}u_{N}v-i\omega v(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u_{N})\right)\,\mathsf{d}\bm{\mathsf{x}}=\lambda_{N}\int_{\Omega}u_{N}v\,\mathsf{d}\bm{\mathsf{x}}\qquad\forall v\in\mathbb{X}_{N},$ (43)
@@
3.2. Fully discrete $hp$–version GFI
We now turn to the spatial discretisation of the forward Euler time-stepping gradient flow iteration (38) on the $hp$–finite element subspace $\mathbb{X}_{N}\subset\mathbb{H}$. To this end, for any $u\in\mathbb{X}_{N}$, we denote by $\mathsf{G}_{N}(u)\in\mathbb{X}_{N}$ the unique solution of
@@ 3.3. Competitive $hp$-refinements and local energy decays
In this section we pave the way for an $hp$–adaptivity algorithm based on employing a competitive refinement strategy on each element $\kappa$ in the computational mesh $\mathcal{T}_{N}$, $N\geq 1$. The essential idea is to compute the maximal predicted energy reduction on each element $\kappa\in\mathcal{T}_{N}$, subject to either a local $h$– or $p$–refinement. Elements with the largest maximal predicted decrease in the local energy functional are then appropriately refined. To this end, we proceed with the following steps.
@@ 3.3.1. Local patch refinements
3.3.2. Local $h$– and $p$–refinements
We now outline the proposed competitive $h$–/$p$–refinements on a given finite element space $\mathbb{X}_{N}=\mathbb{V}(\mathcal{T}_{N},\bm{\mathsf{p}}_{N}$). To this end, for an element $\kappa\in\mathcal{T}_{N}$, we construct the local mesh patches $\mathcal{T}^{\kappa,{\tt ref}}_{N}$ and $\mathcal{T}^{\kappa}_{N}$, as outlined above. Then, local space enrichments in terms of pure $h$–, respectively, $p$–refinement, may be considered as follows:
@@ 3.3.2. Local $h$– and $p$–refinements
    Secondly, we consider the locally refined space $\widehat{\mathbb{V}}^{\kappa}_{N,\rm p}:=\mathbb{V}(\mathcal{T}^{\kappa}_{N},\bm{\mathsf{p}}^{\kappa}_{\rm p})$ for the element $\kappa$ that is solely based on $p$–refinement, where we define a modified (global) polynomial degree vector on $\mathcal{T}_{N}$ by
@@ 3.3.3. Local energy decays
On a given mesh $\mathcal{T}_{N}$, $N\geq 1$, suppose that we have obtained a sufficiently accurate approximation $u_{N}^{n}\in\mathbb{X}_{N}$ of the fully discrete GPE (43), after performing $n\geq 1$ steps of the discrete iteration (45). We now propose an $hp$-type refinement of the mesh $\mathcal{T}_{N}$ based on a competitive local reduction strategy of the energy associated to the available approximation $u_{N}^{n}$. To this end, for each element $\kappa\in\mathcal{T}_{N}$, we collect the associated spaces for the proposed local $h$– and $p$–refinements, cf. above, in the set
@@
where the number of basis function $m^{\kappa}$ depends on the underlying refinement. Then, for the proposed local $h$– or $p$–refinements, and for $u_{N}^{n}\in\mathbb{X}_{N}$ as above, define the augmented space
@@
which allows us to enrich the approximate solution $u_{N}^{n}$ by some local features on $\kappa$ (in terms of the respective $h$– and $p$–refinements). Specifically, we perform one local fully discrete GFI-step (45) in $\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})\subset\mathbb{H}$ in order to obtain a new local approximation, denoted by $\widetilde{u}_{N}^{\kappa,n}\in\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})$, with the usual constraint $\left\|\widetilde{u}_{N}^{\kappa,n}\right\|_{\mathrm{L}^{2}(\Omega)}=1$.
@@
We emphasise that the augmented spaces $\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})$, although they feature one global degree of freedom due to the presence of the given approximation $u_{N}^{n}$ as a basis function, have small dimensions. Hence, in particular, the fully discrete GFI (45) based on $\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})$ entails hardly any computational cost and can be undertaken in parallel for each of the proposed $h$– and $p$–refinements (and for each element $\kappa$).
@@
to be the potential energy reduction, per degree of freedom, when $h$–refinement is employed on $\kappa$, where ${\rm dofs}_{\rm h}^{\kappa}$ denotes the number of degrees of freedom present in the locally $h$–refined space restricted to $\kappa$. Similarly, the potential energy reduction, per degree of freedom, when $p$–refinement is utilised is given by
@@
where ${\rm dofs}_{\rm p}^{\kappa}$ denotes the number of degrees of freedom present in the locally $p$–refined space restricted to $\kappa$. In this way, for each element $\kappa\in\mathcal{T}_{N}$, we may compute the value
@@
where $\theta\in(0,1)$ is a fixed marking parameter. Following [21, 36], cf. also [31], we set $\theta=\nicefrac{{1}}{{3}}$. Once the elements in the current mesh $\mathcal{T}_{N}$ have been $hp$–refined, an $hp$–mesh smoother is then employed in order to construct the enriched finite element space $\mathbb{X}_{N+1}=\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1}$). This is a two-stage process: firstly, hanging nodes created through mesh refinement are removed; here, elements containing two or more hanging nodes are isotropically red refined. This process is repeated until all elements possess at most one hanging node; the remaining hanging nodes are then removed by introducing temporary (green) refinements to yield a conforming mesh $\mathcal{T}_{N+1}$. Subelements created through this process automatically inherit the polynomial degree of their parent. In the second step, the polynomial degree vector is smoothed by ensuring that the maximal difference in the local polynomial degree between adjacent elements is at most one; this is achieved by enriching the element with the lowest order. The above competitive $hp$–refinement strategy is summarised in Algorithm 1.
@@
3:      // $h$–refinement of $\kappa$
4:      Find a basis $\{\xi^{1}_{\kappa},\ldots,\xi^{m^{\kappa}}_{\kappa}\}$ of the locally $h$–enriched space $\widehat{\mathbb{V}}^{\kappa}_{N,\rm h}$.
@@
6:      // $p$–refinement of $\kappa$
7:      Find a basis $\{\xi^{1}_{\kappa},\ldots,\xi^{\widetilde{m}^{\kappa}}_{\kappa}\}$ of the locally $p$–enriched space $\widehat{\mathbb{V}}^{\kappa}_{N,\rm p}$.
@@
12:Perform $h$– or $p$–refinement on each $\kappa\in\mathcal{K}$ according to which refinement takes the maximum in (49) and undertake $hp$-mesh smoothing. This results in a refined global finite element space $\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1})$.
Algorithm 1 Energy-based competitive $hp$-adaptive refinement procedure
3.4. $hp$–adaptive strategy
From a practical viewpoint, once the fully discrete GFI approximation from (45) is close to a solution of the discrete GPE (43), on a given $hp$–finite element space $\mathbb{X}_{N}=\mathbb{V}(\mathcal{T}_{N},\bm{\mathsf{p}}_{N})$, $N\geq 1$, we expect that any further GFI steps will no longer contribute an essential decay of the underlying energy. In this case, in order to further reduce the energy, we need to enrich the finite element space appropriately. Then, the fully discrete GFI iteration is reinitiated on the new space, $\mathbb{X}_{N+1}=\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1})$, and so on. More specifically, for $N\geq 1$, suppose that we have performed a reasonable number $n\geq 1$ (possibly depending on $N$) of GFI-iterations (45) in $\mathbb{X}_{N}$. Consider now an $hp$–enriched space $\mathbb{X}_{N+1}=\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1})$ as obtained by Algorithm 1. Then we may embed (or project, if refinements are not hierarchical) the final guess $u^{n}_{N}\in\mathbb{X}_{N}$ on the previous space into the enriched finite element space $\mathbb{X}_{N+1}$ in order to obtain an initial guess on
@@ 3.4. $hp$–adaptive strategy
for some parameter $10<\gamma<$ 1. We implement this procedure in Algorithm 2.
Algorithm 2 $hp$–Adaptive fully discrete finite element gradient flow iteration
@@
Mark and adaptively $hp$refine the mesh $\mathcal{T}_{N}$ using Algorithm 1 in order to generate a new mesh $\mathcal{T}_{N+1}$.
@@ 4. Numerical Experiments
In this section, we present a series of numerical experiments to investigate the practical performance of the proposed $hp$–refinement algorithm outlined in Algorithms 1 & 2. Throughout this section, we set $\theta=\nicefrac{{1}}{{3}}$, cf. above, and $\gamma=10^{-3}$. The selection of $\mathtt{tol}$ is problem dependent in order to compute a highly accurate value of the ground-state energy for the example at hand. Throughout this section we compare the performance of the proposed $hp$–adaptive refinement strategy with the corresponding algorithm based on exploiting only local mesh subdivision, i.e., $h$–refinement, cf. [25]. Furthermore, to compute the solution of the underlying linear problem (44), we employ the direct MUltifrontal Massively Parallel Solver (MUMPS), see [5, 4, 6].
@@ 4.1. GPE without angular momentum operator $\omega=0$
Firstly, we consider a series of numerical examples where the angular momentum operator is omitted, i.e., when $\omega=0$. In particular, we demonstrate the performance of the proposed $hp$–adaptive refinement algorithm on a series of examples presented in [25]. Throughout this section, we select the initial ground-state $u_{0}^{0}\in\mathbb{X}_{0}\subset\mathbb{H}$, such that $u_{0}^{0}(\bm{\mathsf{x}})=c$ for any node $\bm{\mathsf{x}}$ in the interior of the initial mesh $\mathcal{T}_{0}$, where, $c>0$ is the appropriate real constant which fulfils the norm constraint $\left\|u_{0}^{0}\right\|_{\mathrm{L}^{2}(\Omega)}=1$.
@@ 4.1.1. Non-linear GPE with harmonic confinement potential
In this first example we consider a non-linear Bose-Einstein condensate with $\beta=1000$, $V=\nicefrac{{1}}{{2}}(x^{2}+y^{2})$, ($\omega=0$), on the computational domain $\Omega=(-6,6)^{2}$. This example has been previously considered in [28] and [25], where the approximations $\mathsf{E}(u_{\mathrm{GS}})\approx 11.9860647$ and $\mathsf{E}(u_{\mathrm{GS}})\approx 11.98605121$ have been computed, respectively, for the energy of the ground-state. Based on employing the $hp$–refinement algorithm proposed in this article, cf. Algorithms 1 & 2, we compute a reference value of $\mathsf{E}(u_{\mathrm{GS}})\approx\underline{11.98605114667}1144$. Here, the underlined digits are stable in the sense that our computations indicate that they do not change as further iterations of the $hp$–refinement strategy are computed.
@@
Figure 2. Non-linear GPE with harmonic confinement potential. (a) Computed ground-state $u_{N}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final $hp$-mesh.
The computed approximation to the ground-state $u^{n}_{N}\equiv u_{N}\approx u_{\mathrm{GS}}$ is depicted in Figure 2(a), the shape of which resembles an ‘upside down bowl’. Furthermore, in Figure 2(b) we illustrate the performance of the proposed $hp$–adaptive algorithm, cf. Algorithms 1 & 2. Here, we plot the error of the energy $\mathsf{E}(u_{N})-\mathsf{E}(u_{\mathrm{GS}})$ against the square root of the number of degrees of freedom employed within the finite element space $\mathbb{X}_{N}$, on a linear–log scale, based on employing both $h$– and $hp$–refinement. We observe that the $hp$–refinement algorithm leads to an exponential decay in the error of the approximated ground-state energy as the finite element space $\mathbb{X}_{N}$ is adaptively enriched: indeed, on a linear-log plot, the convergence line is roughly straight. Moreover, we observe the superiority of $hp$–refinement in comparison with a standard $h$–refinement algorithm, cf. [25], in the sense that the former refinement strategy leads to several orders of magnitude reduction in the error in $\mathsf{E}$, for a given number of degrees of freedom, than the corresponding quantity computed exploiting mesh subdivision only.
Finally, in Figure 2(c) we show the final $hp$–refined mesh generated by the proposed refinement strategy. Here we observe that some $h$–refinement has been undertaken around the base of the (upside down) bowl profile of the ground-state, as well as in the centre of the domain. The polynomial degree has then been enriched where the ground-state has steep gradients, as we would expect.

### v2

@@ 3.1. $hp$-finite element discretisation
For the purpose of a computational realisation of the discrete GFI (32), Galerkin spaces $\mathbb{X}_{N}$ will be constructed in terms of an hphp–finite element approach. We focus on a sequence of conforming and shape-regular partitions $\{\mathcal{T}_{N}\}_{N\in\mathbb{N}}$ of the domain $\Omega$ into simplicial elements $\mathcal{T}_{N}=\{\kappa\}_{\kappa\in\mathcal{T}_{N}}$ (i.e., triangles for $d=2$ and tetrahedra for $d=3$), although more general elements could also be considered. For $N\geq 0$ and an element $\kappa\in\mathcal{T}_{N}$, we associate a polynomial degree $p_{\kappa}\geq 1$ and collect these quantities in the polynomial degree vector $\bm{\mathsf{p}}_{N}=[p_{\kappa}:\kappa\in\mathcal{T}_{N}]$. With this notation, for any subset $\varpi\subset\mathcal{T}_{N}$, we introduce the following hphp–finite element space:
@@
$\begin{split}\int_{\Omega}\big(\nabla u_{N}\cdot\nabla\overline{v}+2V(\bm{\mathsf{x}})u_{N}\overline{v}+2\beta|u_{N}|^{2}u_{N}\overline{v}&-2i\omega\overline{v}(\bm{\mathsf{A}}(\bm{\mathsf{x}})\cdot\nabla u_{N})\big)\,\mathsf{d}\bm{\mathsf{x}}\\ &=\lambda_{N}\int_{\Omega}u_{N}\overline{v}\,\mathsf{d}\bm{\mathsf{x}}\qquad\forall v\in\mathbb{X}_{N},\end{split}$ (47)
@@
3.2. Fully discrete hphp–version GFI
We now turn to the spatial discretisation of the forward Euler time-stepping gradient flow iteration (32) on the hphp–finite element subspace $\mathbb{X}_{N}\subset\mathbb{H}$. To this end, for any $u\in\mathbb{X}_{N}$, we denote by $\mathsf{G}_{N}(u)\in\mathbb{X}_{N}$ the unique solution of
@@ 3.3. Competitive $hp$-refinements and local energy decays
In this section we pave the way for an hphp–adaptive algorithm based on employing a competitive refinement strategy on each element $\kappa$ in the computational mesh $\mathcal{T}_{N}$, $N\geq 0$. The essential idea is to compute the maximal predicted energy reduction on each element $\kappa\in\mathcal{T}_{N}$, subject to either a local hh– or pp–refinement. Elements with the largest maximal predicted decrease in the local energy functional are then appropriately refined. To this end, we proceed with the following steps.
@@ 3.3.1. Local patch refinements
3.3.2. Local hh– and pp–refinements
We now outline the proposed competitive hh–/pp–refinements on a given finite element space $\mathbb{X}_{N}=\mathbb{V}(\mathcal{T}_{N},\bm{\mathsf{p}}_{N}$). To this end, for an element $\kappa\in\mathcal{T}_{N}$, we construct the local mesh patches $\mathcal{T}^{\kappa,{\tt ref}}_{N}$ and $\mathcal{T}^{\kappa}_{N}$, as outlined above. Then, local space enrichments in terms of pure hh–, respectively, pp–refinement, may be considered as follows:
@@ 3.3.2. Local hh– and pp–refinements
    Secondly, we consider the locally refined space $\widehat{\mathbb{V}}^{\kappa}_{N,\rm p}:=\mathbb{V}(\mathcal{T}^{\kappa}_{N},\bm{\mathsf{p}}^{\kappa}_{\rm p})$ for the element $\kappa$ that is solely based on pp–refinement, where we define a modified (global) polynomial degree vector on $\mathcal{T}_{N}$ by
@@ 3.3.3. Local energy decays
On a given mesh $\mathcal{T}_{N}$, $N\geq 0$, suppose that we have obtained a sufficiently accurate approximation $u_{N}^{n}\in\mathbb{X}_{N}$ of the fully discrete GPE (47), after performing $n\geq 1$ steps of the discrete iteration (49). We now propose an $hp$-type refinement of the mesh $\mathcal{T}_{N}$ based on a competitive local reduction strategy of the energy associated to the available approximation $u_{N}^{n}$. To this end, for each element $\kappa\in\mathcal{T}_{N}$, we collect the associated spaces for the proposed local hh– and pp–refinements, cf. above, in the set
@@
where the number of basis function $m^{\kappa}$ depends on the underlying refinement. Then, for the proposed local hh– or pp–refinements, and for $u_{N}^{n}\in\mathbb{X}_{N}$ as above, define the augmented space
@@
which allows us to enrich the approximate solution $u_{N}^{n}$ by some local features on $\kappa$ (in terms of the respective hh– and pp–refinements). Specifically, we perform one local fully discrete GFI-step (49) in $\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})\subset\mathbb{H}$ in order to obtain a new local approximation, denoted by $\widetilde{u}_{N}^{\kappa,n}\in\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})$, with the usual constraint $\left\|\widetilde{u}_{N}^{\kappa,n}\right\|_{\mathrm{L}^{2}(\Omega)}=1$.
@@
We emphasise that the augmented spaces $\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})$, although they feature one global degree of freedom due to the presence of the given approximation $u_{N}^{n}$ as a basis function, have small dimensions. Hence, in particular, the fully discrete GFI (49) based on $\widehat{\mathbb{W}}^{\kappa}_{N}(u_{N}^{n})$ entails hardly any computational cost and can be undertaken in parallel for each of the proposed hh– and pp–refinements (and for each element $\kappa$).
@@
to be the potential energy reduction, per degree of freedom, when hh–refinement is employed on $\kappa$, where ${\rm dofs}_{\rm h}^{\kappa}$ denotes the number of degrees of freedom present in the locally hh–refined space restricted to $\kappa$. Similarly, the potential energy reduction, per degree of freedom, when pp–refinement is utilised is given by
@@
where ${\rm dofs}_{\rm p}^{\kappa}$ denotes the number of degrees of freedom present in the locally pp–refined space restricted to $\kappa$. In this way, for each element $\kappa\in\mathcal{T}_{N}$, we may compute the value
@@
where $\theta\in(0,1)$ is a fixed marking parameter. Following [20, 36], cf. also [32], we set $\theta=\nicefrac{{1}}{{3}}$. Once the elements in the current mesh $\mathcal{T}_{N}$ have been hphp–refined, an hphp–mesh smoother is then employed in order to construct the enriched finite element space $\mathbb{X}_{N+1}=\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1}$). This is a two-stage process: firstly, hanging nodes created through mesh refinement are removed; here, elements containing two or more hanging nodes are isotropically red refined. This process is repeated until all elements possess at most one hanging node; the remaining hanging nodes are then removed by introducing temporary (green) refinements to yield a conforming mesh $\mathcal{T}_{N+1}$. Subelements created through this process automatically inherit the polynomial degree of their parent. In the second step, the polynomial degree vector is smoothed by ensuring that the maximal difference in the local polynomial degree between adjacent elements is at most one; this is achieved by enriching the element with the lowest order. The above competitive hphp–refinement strategy is summarised in Algorithm 1.
@@
3:   // hh–refinement of $\kappa$
4:   Find a basis $\{\xi^{1}_{\kappa},\ldots,\xi^{m^{\kappa}}_{\kappa}\}$ of the locally hh–enriched space $\widehat{\mathbb{V}}^{\kappa}_{N,\rm h}$.
@@
6:   // pp–refinement of $\kappa$
7:   Find a basis $\{\xi^{1}_{\kappa},\ldots,\xi^{\widetilde{m}^{\kappa}}_{\kappa}\}$ of the locally pp–enriched space $\widehat{\mathbb{V}}^{\kappa}_{N,\rm p}$.
@@
12:Perform hh– or pp–refinement on each $\kappa\in\mathcal{K}$ according to which refinement takes the maximum in (53) and undertake hphp-mesh smoothing. This results in a refined global finite element space $\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1})$.
Algorithm 1 Energy-based competitive hphp-adaptive refinement procedure
3.4. hphp–adaptive strategy
From a practical viewpoint, once the fully discrete GFI approximation from (49) is close to a solution of the discrete GPE (47), on a given hphp–finite element space $\mathbb{X}_{N}=\mathbb{V}(\mathcal{T}_{N},\bm{\mathsf{p}}_{N})$, $N\geq 0$, we expect that any further GFI steps will no longer contribute an essential decay of the underlying energy. In this case, in order to further reduce the energy, we need to enrich the finite element space appropriately. Then, the fully discrete GFI iteration is reinitiated on the new space, $\mathbb{X}_{N+1}=\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1})$, and so on. More specifically, for $N\geq 0$, suppose that we have performed a reasonable number $n\geq 1$ (possibly depending on $N$) of GFI-iterations (49) in $\mathbb{X}_{N}$. Consider now an hphp–enriched space $\mathbb{X}_{N+1}=\mathbb{V}(\mathcal{T}_{N+1},\bm{\mathsf{p}}_{N+1})$ as obtained by Algorithm 1. Then we may embed (or project, if refinements are not hierarchical) the final guess $u^{n}_{N}\in\mathbb{X}_{N}$ on the previous space into the enriched finite element space $\mathbb{X}_{N+1}$ in order to obtain an initial guess on
@@
for some parameter $0<\gamma<1$. We implement this procedure in Algorithm 2.
Algorithm 2 hphp–Adaptive fully discrete finite element gradient flow iteration
@@
Mark and adaptively hphprefine the mesh $\mathcal{T}_{N}$ using Algorithm 1 in order to generate a new mesh $\mathcal{T}_{N+1}$.
@@ 4. Numerical Experiments
In this section, we present a series of numerical experiments to investigate the practical performance of the proposed hphp–refinement algorithm outlined in Algorithms 1 & 2. Throughout this section, we set $\theta=\nicefrac{{1}}{{3}}$, cf. above, and $\gamma=10^{-3}$. The selection of $\mathtt{tol}$ is problem dependent in order to compute a highly accurate value of the ground-state energy for the example at hand. Throughout this section we compare the performance of the proposed hphp–adaptive refinement strategy with the corresponding algorithm based on exploiting only local mesh subdivision, i.e., hh–refinement, cf. [26]. Furthermore, to compute the solution of the underlying linear problem (48), we employ the direct MUltifrontal Massively Parallel Solver (MUMPS), see [5, 4, 6].
@@ 4.1. GPE without angular momentum operator $\omega=0$
Firstly, we consider a series of numerical examples where the angular momentum operator is omitted, i.e., when $\omega=0$. In particular, we demonstrate the performance of the proposed hphp–adaptive refinement algorithm on a series of examples presented in [26]. Throughout this section, we select the initial ground-state $u_{0}^{0}\in\mathbb{X}_{0}\subset\mathbb{H}$, such that $u_{0}^{0}(\bm{\mathsf{x}})=c$ for any node $\bm{\mathsf{x}}$ in the interior of the initial mesh $\mathcal{T}_{0}$, where, $c>0$ is the appropriate real constant which fulfils the norm constraint $\left\|u_{0}^{0}\right\|_{\mathrm{L}^{2}(\Omega)}=1$.
@@ 4.1.1. Non-linear GPE with harmonic confinement potential
In this first example we consider a non-linear Bose-Einstein condensate with $\beta=1000$, $V=\nicefrac{{1}}{{2}}(x^{2}+y^{2})$, ($\omega=0$), on the computational domain $\Omega=(-6,6)^{2}$. This example has been previously considered in [29] and [26], where the approximations $\mathsf{E}(u_{\mathrm{GS}})\approx 11.9860647$ and $\mathsf{E}(u_{\mathrm{GS}})\approx 11.98605121$ have been computed, respectively, for the energy of the ground-state. Based on employing the hphp–refinement algorithm proposed in this article, cf. Algorithms 1 & 2, we compute a reference value of $\mathsf{E}(u_{\mathrm{GS}})\approx\underline{11.98605114667}1144$. Here, the underlined digits are stable in the sense that our computations indicate that they do not change as further iterations of the hphp–refinement strategy are computed.
@@
Figure 2. Non-linear GPE with harmonic confinement potential. (a) Computed ground-state $u_{N}$; (b) Comparison of the error with respect to the square root of the number of degrees of freedom; (c) Final hphp-mesh.
The computed approximation to the ground-state $u^{n}_{N}\equiv u_{N}\approx u_{\mathrm{GS}}$ is depicted in Figure 2(a), the shape of which resembles an ‘upside down bowl’. Furthermore, in Figure 2(b) we illustrate the performance of the proposed hphp–adaptive algorithm, cf. Algorithms 1 & 2. Here, we plot the error of the energy $\mathsf{E}(u_{N})-\mathsf{E}(u_{\mathrm{GS}})$ against the square root of the number of degrees of freedom employed within the finite element space $\mathbb{X}_{N}$, on a linear–log scale, based on employing both hh– and hphp–refinement. We observe that the hphp–refinement algorithm leads to an exponential decay in the error of the approximated ground-state energy as the finite element space $\mathbb{X}_{N}$ is adaptively enriched: indeed, on a linear-log plot, the convergence line is roughly straight. Moreover, we observe the superiority of hphp–refinement in comparison with a standard hh–refinement algorithm, cf. [26], in the sense that the former refinement strategy leads to several orders of magnitude reduction in the error in $\mathsf{E}$, for a given number of degrees of freedom, than the corresponding quantity computed exploiting mesh subdivision only.
Finally, in Figure 2(c) we show the final hphp–refined mesh generated by the proposed refinement strategy. Here we observe that some hh–refinement has been undertaken around the base of the (upside down) bowl profile of the ground-state, as well as in the centre of the domain. The polynomial degree has then been enriched where the ground-state has steep gradients, as we would expect.